An empty steel container is filled with 0.0500 atm of HF. The system is allowed to reach equilibrium. If Kp = 2.76 for the reaction below, what is the equilibrium partial pressure of H₂?
1
Expert's answer
2019-11-19T07:59:35-0500
"2HF = H_2 + F_2"


Suppose the system initially contained 2 mol HF, then:


"K_p = {x^2 \over 2 - 2x}"




"2.76(2 - 2x) = x^2"

"5.52 - 5.52x = x^2"

"x^2 + 5.52x - 5.52 = 0"

"x = 0.86"

Then, the amount of substance substances in the reaction in equilibrium:


"n(HF) = 0.28 mol"

"n(H_2) = n(F_2) = 0.86 mol"

"\chi(H_2) = {0.86 \over 0.86 + 0.86 + 0.28} = 0.43"

Then:


"P_{H_2} = \chi(H_2)*P"

"P_{H_2} = 0.43*0.0500atm = 0.0215atm"

Answer: 0.0215 atm